Container 1 has 8 items, 3 of which are defective. Container 2 has 5 items, 2 of which are defective. If one item is drawn from each container, what is the probability that both items are not defective? 0.3750 0.3846 0.1500 0.6154 show work please
Answer #1

If one item is drawn from each container then the probability that both the items are not defective

\\\\=\frac{5}{8}*\frac{3}{5} \\\\=\frac{15}{40} \\\\=0.3750

First option is correct.